jhonyy9 (jhonyy9):

than given lg 4 = p rewrite lg 25 in function of p (using p) lg mean logarithm base ten i ve got lg 25 = ((2-p)/p)lg 4 so what proven

jhonyy9 (jhonyy9):

proven will be ((2-p)/p)lg 4 = ((2-p)/lg 4)lg 4 where lg 4 with lg 4 from denominator simplified and will remain there (2-p) so lg 25 = 2 -p lg 25 = 2 -lg 4 lg 25 = 2lg 10 -lg 4 lg 25 = lg 10^2 -lg 4 lg 25 = lg 100 - lg 4 lg 25 = lg (100/4) lg 25 = lg 25 hope this is right sure

jhonyy9 (jhonyy9):

@ganeshie8 @TheSmartOne @mathmate @Kainui your opinion about this please ? ty

jhonyy9 (jhonyy9):

agent0smith how you see it please ?

OpenStudy (agent0smith):

If \[\log 4 = p\]then \[\large \log 25 = \frac{ \log 25 }{ \log 4 }*p\]since p=log 4.

jhonyy9 (jhonyy9):

yes i understand your idea but there on the right side dont use just lg 4 or p because you need rewriting the lg 25 in function of p

jhonyy9 (jhonyy9):

using the properties of logarithms lg 4 = p => 10^p = 4 - yes ? (2^p)*(5^p) = 2^2 5^p = (2^2)/(2^p) 5^p = 2^(2-p) make radical index p from both sides or push both sides on power of 1/p 5 = (2^(2-p))^(1/p) 5 = 2^((2-p)/p) squared both sides 5^2 = 2^(((2-p/p))*2) 5^2 = 4^((2-p/p) lograith lg on both sides lg 25 = ((2-p/p)) lg 4 but we know that p=lg 4 lg 25 = ((2-lg 4)/lg 4)lg 4 lg 25 = (2-lg 4) lg 25 = lg 10^2 -lg 4 lg 25 = lg 100 - lg 4 lg 25 = lg (100/4) lg 25 = lg 25 i think this is right in this way

jhonyy9 (jhonyy9):

@ShadowLegendX @MrNood opinions please ty

jhonyy9 (jhonyy9):

@sweetburger

OpenStudy (agent0smith):

Yeah that looks fine \[\large \log 25 = \frac{ 2-p }{ p } *p\]or log 25 = 2 - p

jhonyy9 (jhonyy9):

thank you and i think this is unbeliveblie but i solved it yestoday in my workplace dont using any from so without pen,paper or calculator and today when i ve said this to my student collage there have said me that i m crazy bc. how is this possible ? but when i ve said him this result of ((2-p)/p)lg 4 said me that is very very nice

OpenStudy (mrnood):

your 'proof' is a trivial exercise (I mean mathematically - not effortwise1)

jhonyy9 (jhonyy9):

ty your opinion Nood

OpenStudy (mrnood):

a=a is self evident I expect your proof is 'circular - i.e. you derived 1 expression and then went 'back the same way

OpenStudy (mrnood):

25 = 100/4 log 25 = log 100 - log 4 log 25 = 2 -p

jhonyy9 (jhonyy9):

ok. but you know a different way to rewrite lg 25 in function of p when you know that p=lg 4 ? wait and thank you

jhonyy9 (jhonyy9):

but please ATTENTION you need starting from lg 4 = p

OpenStudy (agent0smith):

@MrNood's method is easier. I don't buy the "rules" that you need to start from a certain expression.

jhonyy9 (jhonyy9):

ok. thank you

jhonyy9 (jhonyy9):

@zepdrix your opinion please about this ? ty

OpenStudy (mathmate):

log 4 does not equal log 25, so cannot start from log 4. The question is "rewrite lg 25 in function of p", which means we start with log 25. I endorse @MrNood's approach.

jhonyy9 (jhonyy9):

ty. @ganeshie8 your opinion please about these ? ty.

jhonyy9 (jhonyy9):

@pooja195 @Callisto

OpenStudy (mww):

Here's another way similar to Mr. Nood's approach \[\log(4) = \log(2^2) = 2\log(2)\] \[\log(25) = \log(5^2)=2\log(5)\] Adding these we get \[p + \log(25) = 2(\log(2) +\log(5)) = 2\log(10) = 2\] Thus log(25) = 2 - p

jhonyy9 (jhonyy9):

@kl0723 np. welcome - sorry but i not can sending to you message just in this way because my submit button was blocked and not is solved this problem again and i not can sending messages to nobody - what i ve wrote there i see it true on this way - good luck

jhonyy9 (jhonyy9):

@Directrix what is your opinion about this please dear my friend ?

OpenStudy (mrnood):

@jhonyy9 what do you actually WANT from your various 'tagged' friends? the answer is clear above.

OpenStudy (agent0smith):

^exactly. There's no need for more input.